INFRARED RESONANCE ENHANCED PHOTODISSOCIATION SPECTROSCOPY OF Fe^{+}-(CO_{2})_{n} CLUS-TERS.

\underline{G. GREGOIRE}, J. VELASQUEZ, M. A. DUNCAN, Department of Chemistry, University of Georgia, Athens, GA 30602, USA.

Mid-infrared (2100-2900 cm^{−1}) photodissociation spectra of mass selected Fe^{+}-(CO_{2})_{n=2−6} clusters are presented. The observed fragmentation channels correspond to the loss of CO_{2} molecules, their numbers depending on the infrared laser intensity. In low power condition, namely a few mJ/cm^{2}, clusters undergo simple evaporation of one CO_{2} molecule, while at higher power (hundreds of mJ/cm^{2}), multi photon absorption leads to multiple fragment detection, ranking from Fe^{+}-(CO_{2})_{n−1} to Fe^{+}-(CO_{2})_{2}. In the latter case, step-by-step solvent evaporation within the temporal width of the laser (8ns) is observed, reflecting that the dissociation rate constant is faster than a few nanosecond. Photodissociation spectra of these clusters show a general blue shift of the resonant CO_{2} asymmetric stretch within the complexes, and this shift decreases as the cluster size increases, which is consistent with the lowering of the Fe^{+}-CO_{2} interaction at the expense of the solvation. Furthermore, weaker transitions are observed which could correspond to combination bands with low frequency mode of about 20 cm^{−1}.